A heavy block is suspended from a vertical string the elastic potential energy stored in the spring is 4J what is the spring constant if the elongation of the spring is 10cm?


1
Expert's answer
2021-06-11T11:21:35-0400

We can find the spring constant from the definition of the elastic potential energy:


"PE_s=\dfrac{1}{2}kx^2,""k=\dfrac{2PE_s}{x^2},""k=\dfrac{2\cdot4\ J}{(0.1\ m)^2}=800\ \dfrac{N}{m}."